How much work does the electric field do in moving a proton from a point with a potential of +165V to a point where it is -45V ? Express your answer in joules.

Express your answer in electron volts.
Answer #1

work = q(v2 - v1)

           = 1.6*10^-19 * (165 -(-45))

           = 3.36*10^-17 J

             = 210 ev
Answer #2

Work done in an electricfield is

   W   =   charge(q) * potenial difference (V)

   HereV   =   - 45 V - 165V   =   - 210 V

   W   =   1.6* 10-19 * ( -210V)   =    - 3.36 *10-17   J

   -ve sign indicates that workis done by electric field.

   W ineV   =   W in joule / 1.6 *10-19    = 3.36 *10-17 / 1.6 * 10-19

   W   = 210 eV
Answer #3

Given that,

Initial Potentail (Vi) = +165 V

FInal Potentail (Vf) = - 45 V

Work Done by electric feild in moving a proton is to be calculated.

Elecrtic Charge of a proton = Qp = 1.602